Find the length of the missing side of the right triangle. Round to three decimal places, if necessary.
a= 12, c = 20
Answer #1
Find the missing side of a right triangle in which a = 12, c = 20, b = ? Assuming that side c is the hypotenuse, then you can use the Pythagorean theorem to find side b: {{{c^2 = a^2+b^2}}} Substitute a = 12 and c = 20. {{{20^2 = 12^2+b^2}}} Simplify. {{{400 = 144+b^2}}} Subtract 144 from both sides of the equation. {{{256 = b^2}}} Take the square root of both sides. {{{highlight(b = 16)}}}
answered by: Not putting it
Answer #2

IDK, I came one here for answers and I never really knew how to do this

source: MYSELF
answered by: NUNYA